OpenStudy (anonymous):

21. –6p + 7 = 3(2p – 3) – 4(–10 + 4p) (1 point) p = 6 p = 5 p = 7 p = 12

OpenStudy (anonymous):

p=6

OpenStudy (agreene):

p=6 just plug in 6 for p and evaluate, you will get -29 = -29

OpenStudy (anonymous):

-6p + 7 = 6p - 9 + 40 - 16p -6p + 7 + 6p = 6p - 9 + 40 - 16p + 6p 7 = 6p - 9 + 40 - 16p + 6p 7 = 12p - 9 + 40 - 16p 7 = -4p + 31 -24 = -4p -24/-4 = -4p/-4 6 = p

OpenStudy (anonymous):

That's a real pain to write out. First use the distributive property on the right hand side Then isolate the variables, separating p combine like terms and solve for p (just as sean walsh patiently posted.)